Exercise 2 Find V2 and V3? 200 k12 AAA 20 k 2 w Vi = 0.2 V 200k 10k2
Answer #1

touz YN V + + 1 Vic 0127 TK robook 200k Ilok By virtual ground property the voltage at inverting and non investing terminal aFor opamp ② . -0.2V Ntana so v = Na = 0.20 terminal By KCL at w 0.20 0.2 – V3 70 200 K lok 20 N2 = 0.2 x 20 +0.2 V3 = 4 +0.2